OpenStudy (anonymous):

will give medal A plot of land is rectangular and has an area of x^2-5x-24m^2. The length is x+3m. Find the width.

OpenStudy (mertsj):

Divide x^2-5x-24 by x+3

OpenStudy (anonymous):

yes i understand that part but when you divide it gets confusing!

OpenStudy (mertsj):

Can you factor x^2-5x-24?

OpenStudy (anonymous):

nope

OpenStudy (mertsj):

Think of two numbers. When you add them you get -5. When you multiply them you get -24

OpenStudy (anonymous):

3 and -8?

OpenStudy (mertsj):

So: \[x^2-5x-24=(x-8)(x+3)\]

OpenStudy (anonymous):

oooooooohhhhhhhhh thanks man appreciate it!!

OpenStudy (mertsj):

And: \[\frac{x^2-5x-24}{x+3}=\frac{(x-8)(x+3)}{x+3}\]

OpenStudy (mertsj):

Cancel the common factor.

OpenStudy (anonymous):

hold on where did the m^2?

OpenStudy (mertsj):

I'm pretty sure that m^2 means square meters

OpenStudy (anonymous):

ohhh never mind

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw